OpenStudy (anonymous):

Determine the radius of a circle whose arc length measures 57.1 meters and central angle measures 1.46 radians.

OpenStudy (jdoe0001):

\(\large \bf \textit{arc's length}=s=r\theta\implies \cfrac{s}{\theta}=r \\ \quad \\ r=radius\qquad \theta=angle\ in\ radians\)